# Towards Multi-agent Policy-based Directed Hypergraph Learning for Traffic Signal Control

### v1

@@ III-B1 Directed Hypergraph
We calculate multi-head attention between the master node and the two types of hyperedge ($e_{spa}\left(\dot{v}^{t}_{i}\right),e_{tem}\left(\dot{v}^{t}_{i}\right)$). The calculation process for the weights of the two types of hyperedge is similar, and thus we obtain spatial hyperedge attention by:
$att^{h}\left(\dot{v}^{t}_{i}\right)=\frac{\emph{H}_{\mathcal{G}}^{t}\left(\dot{v}^{t}_{i}\right)\cdot Q\text{-}Lin^{h}\cdot\Theta_{spa}^{att}\cdot\emph{K}_{spa}^{h}\left(\dot{v}^{t}_{i}\right)^{T}}{\sqrt{d}}$ (7)
@@
Here, for the $h$-th attention head $att^{h}\left(\dot{v}^{t}_{i}\right)$, $Q\text{-}Lin^{h}\in\mathcal{R}^{d\times\frac{d}{K}}$ is linear transformation matrix, and $K\text{-}Lin^{h}_{spa}$ is a trainable matrix with the same dimensions. $K$ is the number of attention heads. $\Theta_{spa}^{att}\in\mathcal{R}^{\frac{d}{K}\times\frac{d}{K}}$ is trainable weight matrix. The $h$-th temporal attention $att^{h}\left(\dot{v}^{t}_{i}\right)$ is calculated in a similar manner. The weight of the spatial hyperedge $w^{h}_{spa}\left(\dot{v}^{t}_{i}\right)$ and the temporal hyperedge $w^{h}_{tem}\left(\dot{v}^{t}_{i}\right)$ are calculated by $softmax$ normalization to $att^{h}\left(\dot{v}^{t}_{i}\right)$ and $att^{h}\left(\dot{v}^{t}_{i}\right)$.
@@
III-B2 Multi-agent Deep Reinforcement Learning
PPO-Clip[30] modifies the surrogate objective by clipping the probability ratio, to penalize changes to the policy that move $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}$ away from 1. The policy loss function is defined as:
@@ III-B2 Multi-agent Deep Reinforcement Learning
where $\epsilon$ is a hyperparameter indicating the range for clipping. $\hat{A}_{t}$ is an estimator of the advantage function at time $t$ [30].
To realize the procedure in an end-to-end fashion, the loss function used in the training process is denoted as:
@@
IV Experiments
We establish a simulation testbed based on CityFlow [32], which is an open-source traffic simulator that supports large-scale traffic signal control. Two real-world datasets $D_{\textit{Hangzhou}}$ and $D_{\textit{Jinan}}$ [13] are adopted in the experiment.
We introduce the following five methods as baseline: Fixed-time [19]: Assigning a constant cycle length with a predetermined green ratio distribution across all phases. MaxPressure [21]: Controlling traffic signals by alleviating congestion on the upstream and downstream queue length with the max pressure. GCN [12]: Using graph convolution neural network to extract traffic features among multiple intersections. PressLight [24]: Integrating pressure into the state and reward design of the reinforcement learning model to solve the multi-intersection signal control problems. MPLight [25]: A decentralized network level traffic signal control RL algorithm with parameter sharing which enables large scale application based on Presslight. CoLight [13]: Using a graph attention network to enhance coordination and decision-making of multi-intersection traffic signal control.
TABLE I: Performance of average travel time and throughput
  Method Average travel time(s) Throughput
$D_{\textit{Hangzhou}}$ $D_{\textit{Jinan}}$ $D_{\textit{Hangzhou}}$ $D_{\textit{Jinan}}$
  Fixed-time 718.29 814.11 1989 3475
MaxPressure 407.17 343.90 2664 5613
  GCN 709.70 596.67 2760 6024
PressLight 416.73 311.39 2832 6116
MPLight 319.71 304.93 2927 6125
CoLight 302.37 309.93 2930 6116
DHG-PPO 292.56 277.94 2930 6158
Table I summarizes the performance of our proposed DHG-PPO against baselines in real-world datasets. The results demonstrate that DHG-PPO outperforms the baselines in minimizing average travel time (ATT) and maximizing throughput metrics. Here, ATT is defined as the average time spent by all vehicles between entering and leaving the area, while throughput is defined as the number of trips completed by vehicles throughout the simulation.
For $D_{\textit{Jinan}}$, DHG-PPO achieves a maximum reduction in ATT of 8.85% compared to the second optimal solution MPLight. Additionally, the maximum enhancement of throughput over the second-best MPLight reaches 0.53%. Traditional control methods (Fixed-time and MaxPressure) rely on predetermined strategies to regulate traffic signals, which results in unsatisfactory performance in terms of ATT and throughput. In contrast, GCN and CoLight utilize GNNs to enable interactions between multiple intersections to improve overall efficiency. However, in real-world datasets, the dynamic changes in traffic flow create higher-order correlations between intersections that traditional graphs cannot fully capture. PressLight and MPLight incorporate the concept of Maxpressure at each intersection but do not account for the dynamic changes in traffic flow that introduce complex directed spatio-temporal attributes to the entire road network.
@@ IV Experiments
Figure 2: Ablation results on two real-world datasets.
Figure 2 plot the results of albation study. Here, MA-PPO utilizes a different graph module (GAT) from our proposed model to capture correlations between intersections. In contrast, MA-SAC and DHG-SAC incorporate the SAC algorithm to validate the effectiveness of hypergraphs. The results show that DHG-PPO and DHG-SAC converge more rapidly and smoothly, which validates our motivation that hypergraphs offer superior performance in capturing correlations between multiple sets of intersections.
Figure 3 investigates the impact of the hyperparameter $\beta$ in Equation 11. The results indicate that the optimal values of $\beta$ are 0.01 and 0.3 for the $D_{\textit{Hangzhou}}$ and $D_{\textit{Jinan}}$, respectively. As the value of $\beta$ increases, the proportion of the loss associated with dynamically constructing directed hypergraphs rises, which deteriorates the performance of ATT.
@@
Figure 3: Impact of hypergraph loss coefficient $\beta$.
V Conclusion
@@
In this paper, we propose a multi-agent reinforcement learning model DHG-PPO, which incorporates a directed hypergraph module to capture the correlation in road traffic by dynamically constructing spatio-temporal hyperedges. Extensive experimental results demonstrate the superiority of DHG-PPO against the baselines in traffic signal control. Our future work is to enhance the robustness of our method by studying the data cleaning schemes to mitigate the impact of missing or noisy existing in the collected data.

### v2

@@ 4.3 Directed Hypergraph Learning
where $\mathcal{H}_{H}$ and $\mathcal{H}_{T}$ denote the set of head nodes and the set of tail nodes, respectively.
@@
We calculate multi-head attention between the master node and the two types of hyperedge ($e_{spa}\left(\dot{v}^{t}_{i}\right),e_{tem}\left(\dot{v}^{t}_{i}\right)$). The calculation process for the weights of the two types of hyperedge is similar, and thus we obtain spatial hyperedge attention $att^{h}_{spa}\left(\dot{v}^{t}_{i}\right)$ by:
$att^{h}_{spa}\left(\dot{v}^{t}_{i}\right)=\frac{\emph{H}_{\mathcal{G}}^{t}\left(\dot{v}^{t}_{i}\right)\cdot Q\text{-}Lin^{h}\cdot\Theta_{spa}^{att}\cdot\emph{K}_{spa}^{h}\left(\dot{v}^{t}_{i}\right)^{T}}{\sqrt{d}}$ (8)
@@
Here, for the $h$-th attention head $att^{h}\left(\dot{v}^{t}_{i}\right)$, $Q\text{-}Lin^{h}\in\mathcal{R}^{d\times\frac{d}{K}}$ is the linear transformation matrix, and $K\text{-}Lin^{h}_{spa}$ is a trainable matrix with the same dimensions. $K$ is the number of attention heads. $\Theta_{spa}^{att}\in\mathcal{R}^{\frac{d}{K}\times\frac{d}{K}}$ is trainable weight matrix. The $h$-th temporal attention $att^{h}_{tem}\left(\dot{v}^{t}_{i}\right)$ is calculated in a similar manner. The weight of the spatial hyperedge $w^{h}_{spa}\left(\dot{v}^{t}_{i}\right)$ and the temporal hyperedge $w^{h}_{tem}\left(\dot{v}^{t}_{i}\right)$ are calculated by $softmax$ normalization to $att^{h}_{spa}\left(\dot{v}^{t}_{i}\right)$ and $att^{h}_{tem}\left(\dot{v}^{t}_{i}\right)$.
@@
4.4 Multi-agent Deep Reinforcement Learning
PPO-Clip [20] modifies the surrogate objective by clipping the probability ratio, to penalize changes to the policy that move the value of $r_{t}(\theta)$ away from 1. $r_{t}(\theta)$ is defined as:
$r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}$ (11)
The policy loss function is obtained by:
@@ 4.4 Multi-agent Deep Reinforcement Learning
where $\epsilon$ is a hyperparameter indicating the range for clipping. $\hat{A}_{t}$ is an estimator of the advantage function at time $t$:
$\hat{A}_{t}=\delta_{t}+(\gamma\lambda_{GAE})\delta_{t+1}+\cdots+(\gamma\lambda_{GAE})^{T-t+1}\delta_{T-1}$ (13)
where the discount rate $\gamma$ is used to balance short-term and long-term rewards. $\lambda_{GAE}$ is the smoothing coefficient. The temporal difference error is calculated by:
$\delta_{t}=r_{t}+\gamma V(s_{t+1})-V(s_{t})$ (14)
To realize the procedure in an end-to-end fashion, we design the loss function used in the training process as:
@@
4.5 Time Complexity Analysis
The observation embedding layer is involved with time complexity $O(Nd^{2})$, where $N$ is the number of nodes, $d$ is the feature dimension. For the actor network, the time complexity is $O(Nd^{2})$. In critic network, the time complexity of directed hypergraph construction is $O(Nd^{2}+N^{2}d)$. The time complexity of directed hypergraph learning is $O(Nd+NN+Nd^{2}+N^{2}d)\approx O(Nd^{2}+N^{2}d)$. The time complexity of graph readout is $O(Nd^{2})$. Hence, the overall time complexity is $O(6Nd^{2}+2N^{2}d+Nd+NN)\approx O(Nd^{2}+N^{2}d)$.
5 Experiments
5.1 Experimental Setup
We establish a simulation testbed based on CityFlow [36], which is an open-source traffic simulator that supports large-scale traffic signal control. After feeding the given traffic datasets into the simulator, vehicles navigate towards their destinations according to the defined environmental parameters. This simulator assesses the road conditions based on the specified traffic signal control strategy and models the actions of each vehicle, offering the details of the traffic evolution within the road network. The main parameter configurations for our proposal are presented in Table 1.
5.2 Dataset
We introduce four datasets (Table 2) to evaluate the effectiveness of our proposed method: two synthetic datasets, $D_{\textit{4$\times$4}}$ and $D_{\textit{6$\times$6}}$, along with two real-world datasets, $D_{\textit{Hangzhou}}$ and $D_{\textit{Jinan}}$. The specifics of these datasets are outlined below.
  • Synthetic Datasets: We examine two synthetic traffic datasets of varying scales. For the $D_{\textit{4$\times$4}}$ dataset, the simulator generated a total of 1,473 traffic flow data records following the Gaussian distribution. All vehicles entered and exited the network from the edge roads on the periphery. The turning ratios of the vehicles were 10% for left turns, 60% for going straight, and 30% for right turns. In the $D_{\textit{6$\times$6}}$ dataset, vehicles arrived uniformly, with a flow of 300 vehicles per lane per hour in the west-east direction and 90 vehicles per lane per hour in the south-north direction.
  • Real-world Datasets: Two real-world datasets $D_{\textit{Hangzhou}}$ and $D_{\textit{Jinan}}$ [31] are adopted in the experiment. Specifically, the road layout in $D_{\textit{Hangzhou}}$ consists of a $4\times 4$ grid located in Gudang Sub-district. In a similar manner, $D_{\textit{Jinan}}$ features a $3\times 4$ grid in Dongfeng Sub-district of Jinan. Both datasets provide information regarding the road network configurations and the associated traffic flow details. The distribution of intersections within these two datasets is illustrated in Figure 2.
@@ 5.2 Dataset
Figure 2: Road networks for real-world datasets.
Table 1: Main parameter settings.
Parameter Value
Batch size 50
Episodes 100
Learning rate 0.0003 (actor & critic)
Replay memory ($\textit{Thres}_{\textit{size}}$) 1000
Optimizer Adam
Head number (K) 1
Discount rate ($\gamma$) 0.91
Smoothing coefficient ($\lambda_{GAE}$) 0.86
Reconstruction error ($\lambda$) 0.001
Regularizing factor ($\gamma_{2}$) 0.2
Clipping parameter ($\epsilon$) 0.3
Table 2: Traffic flow details of different traffic datasets.
Dataset Intersections Total Vehicle Flow Type
$D_{\textit{Hangzhou}}$ 16 6984 Real-world
$D_{\textit{Jinan}}$ 12 6295 Real-world
$D_{\textit{4$\times$4}}$ 16 1473 Synthetic
$D_{\textit{6$\times$6}}$ 36 3000 Synthetic
5.3 Baselines
We use twelve methods as baselines to validate the effectiveness of our proposed DHLight. The baselines encompass traditional transportation methods, reinforcement learning techniques, and those that integrate graph learning. To ensure a fair comparison, we utilize the optimal parameters configurations for various datasets as specified in the original open source codes of the baseline methods. The details of the baseline methods are as below:
  • Fixed-time [11]: Assigning a constant cycle length with a predetermined green ratio distribution across all phases.
  • MaxPressure [23]: Controlling traffic signals by alleviating congestion on the upstream and downstream queue length with the max pressure.
  • GCN [16]: Using graph convolution neural network to extract traffic features among multiple intersections.
  • PressLight [30]: Integrating pressure into the state and reward design of the reinforcement learning model to solve the multi-intersection signal control problems.
  • CoLight [31]: Using a graph attention network to enhance coordination and decision-making of multi-intersection traffic signal control.
  • MPLight [5]: A decentralized network level traffic signal control RL algorithm with parameter sharing which enables large scale application based on Presslight.
  • Attn-CommNet [8]: A reinforcement learning method that incorporates a local attention mechanism, leveraging local selection and attention between hidden layers to enhance cooperation among agents.
  • GPLight [13]: A multi-agent reinforcement learning method that maintains a balance between accuracy and complexity in large-scale traffic signal control systems by clustering agents that share a high degree of similarity.
  • SO2 [39]: An offline-to-online reinforcement learning method that improves Q-value estimation by perturbing the target action and boosting the frequency of Q-value updates.
  • EMC [26]: A multi-agent coordination method that models each intersection as an autonomous agent, using a cost function to optimize traffic signal decisions and minimize vehicle travel time through decentralized message passing.
  • CrossLight [22]: It combines offline meta-training with online adaptation to facilitate the transfer of traffic signal control strategies between different cities.
  • MixLight [33]: A multi-agent reinforcement learning approach that employs an executor-guide dual network to facilitate bilateral cooperation in multi-intersection TSCSs.
Table 3: The ATT performance of different methods.
Method [Source] Average travel time (secs)
$D_{\textit{Hangzhou}}$ $D_{\textit{Jinan}}$ $D_{\textit{4$\times$4}}$ $D_{\textit{6$\times$6}}$
Fixed-time 718.29 814.11 213.04 334.68
MaxPressure 407.17 343.90 159.56 195.49
GCN [ITSC18’] 709.70 596.67 - 272.14
PressLight [KDD19’] 416.73 311.39 232.04 256.89
CoLight [CIKM19’] 302.37 309.93 141.79 173.88
MPLight [AAAI20’] 319.71 304.93 177.01 238.27
Attn-CommNet [ICTAI21’] 384.00 334.30 - 209.30
GPLight [IJCAI23’] 301.45 307.52 - -
SO2 [AAAI24’] 333.29 281.02 135.08 -
EMC [TITS24’] 362.10 304.10 - -
CrossLight [KDD24’] 329.98 279.52 128.73 -
MixLight [TII24’] 317.90 303.04 - 227.99
DHLight (Ours) 292.56 277.94 123.51 166.10
  • -- indicates results are inapplicable to certain dataset.
@@ 5.3 Baselines
(c) $D_{\textit{4$\times$4}}$
(d) $D_{\textit{6$\times$6}}$
Figure 3: Ablation results on different datasets.
(a) $D_{\textit{Hangzhou}}$, $\beta$
(b) $D_{\textit{Jinan}}$, $\beta$
(c) $D_{\textit{Hangzhou}}$, $\zeta$
(d) $D_{\textit{Jinan}}$, $\zeta$
@@
Figure 4: Impact of hypergraph loss coefficient $\beta$ and threshold $\zeta$ on real-world datasets ($D_{\textit{Hangzhou}}$ and $D_{\textit{Jinan}}$).
5.4 Performance Evaluation
In the field of traffic signal control, Average Travel Time (ATT) is commonly used as a metric to evaluate the performance of TSCSs. ATT is defined as the average time all vehicles spend traveling between entering and exiting a designated area. This value is influenced by factors such as red and yellow traffic signals, as well as traffic congestion. A lower ATT indicates better performance of the TSCSs.
Table 3 summarizes the performance of our proposed DHLight against baselines in four datasets. The results demonstrate that DHLight outperforms the baselines in minimizing ATT: Traditional control methods (Fixed-time and MaxPressure) rely on predetermined strategies to regulate traffic signals, which results in unsatisfactory performance in terms of ATT. In contrast, GCN and CoLight utilize GNNs to enable interactions between multiple intersections to improve overall efficiency. However, in real-world datasets, the dynamic changes in traffic flow create higher-order correlations between intersections that traditional graphs cannot fully capture. PressLight and MPLight incorporate the concept of MaxPressure at each intersection but do not account for the dynamic changes in traffic flow that introduce complex spatio-temporal attributes to the entire road network. The Attn-CommNet model incorporates a local attention mechanism to improve cooperation among agents by facilitating local selection and attention between hidden layers. However, this approach may fall short in capturing global information in complex scenarios, limiting the effectiveness of agent collaboration and ultimately affecting the model’s performance. GPlight is effective in balancing accuracy and complexity through agent clustering, but it struggles to capture intricate interactions among agents in diverse traffic scenarios. SO2 enhances Q-value estimation by perturbing the target action and increasing the frequency of Q-value updates; however, the performance of this method is constrained by its inability to fully address the inherited estimation bias and inaccurate Q-value ranking resulting from offline pretraining while applying to TSCSs. EMC models each intersection as an autonomous agent, using a cost function to optimize traffic signal decisions and minimize vehicle travel time through decentralized message passing; however, it is limited in capturing complex higher-order spatio-temporal relationships, failing to fully leverage the intricate interactions among traffic flows. CrossLight facilitates the transfer of traffic signal control strategies between cities through offline meta-training and online adaptation; however, it inadequately models the spatio-temporal traffic differences across cities, which restricts its transferability. Similarly, while MixLight employs an executor-guide dual network to promote bilateral cooperation in multi-agent systems, it fails to sufficiently capture global spatio-temporal dependencies, potentially hindering its performance in large-scale traffic networks.
In contrast, our proposed DHLight effectively addresses the issues present in baseline methods by dynamically constructing directed hyperedges. By forming directed hyperedges that connect multiple intersections rather than being limited to actual geographical connections, DHLight is able to capture and represent the complex directed spatio-temporal attributes inherent in TSCSs. This innovative approach enables a more comprehensive understanding of the interdependencies among various traffic nodes, allowing for deeper insights into how changes in one area can affect others. By elucidating these relationships, DHLight enhances decision-making processes and improves adaptability within dynamic traffic environments. This capability is crucial for effectively responding to real-time traffic conditions, ultimately leading to more efficient traffic management and optimized flow.
5.5 Ablation Study
Figure 3 plots the results of ablation study for our proposed DHLight. “w/o DHG module” replaces the directed hypergraph module with a traditional pairwise GAT to capture correlations between intersections.The results demonstrate that DHLight achieves satisfactory ATT performance more rapidly and smoothly compared to the traditional graph learning approach (“w/o DHG module”). This result can be attributed to the capability of directed hypergraphs to simultaneously represent relationships involving multiple intersections, which is particularly valuable in traffic networks where interactions often involve more than two intersections. By utilizing directed edges that connect multiple nodes, DHLight can model complex traffic patterns more accurately, capturing the nuances of traffic flow between various intersections. The findings demonstrate the exceptional capability of directed hypergraphs in capturing correlations among multiple intersections.
Similarly, “w/o PPO module” replaces the MA-PPO-based method with the ones with the Multi-Agent Soft Actor-Critic (MA-SAC). From Figure 3, it is clear that DHLight achieves lower and more stable ATT during the first half of the experiments, ultimately yielding comparable results by the end. The underlying reason for this phenomenon is that the MA-PPO method provides more stable policy updates during the initial training phase, leading to the development of higher-quality initial policies. This stability allows the model to quickly and effectively capture the dynamic characteristics of the traffic network, facilitating a more responsive adaptation to changing conditions. In contrast, while the MA-SAC method boasts stronger exploration capabilities, its lower quality of initial policy adversely impacts ATT performance in the early stages of training. The exploratory nature of MA-SAC can lead to suboptimal decision-making as it navigates uncharted areas of the policy space, hindering early performance compared to the more stable and effective approach of MA-PPO. Overall, the advantages of the MA-PPO algorithm lies in its ability to formulate an effective initial policy, which is crucial for efficiently addressing complex traffic dynamics.
5.6 Sensitivity Analysis
Figure 4 investigates the impact of the hyperparameter $\beta$ and $\zeta$ in Equation 15 and Equation 4, respectively. $\beta$ is the weight coefficient assigned to the hyperedge construction losses within the model. This coefficient plays a pivotal role in balancing the contributions of different loss components during the training process, thereby influencing the overall effectiveness of hyperedge formation. The findings reveal that the optimal value of $\beta$ for both $D_{\textit{Hangzhou}}$ and $D_{\textit{Jinan}}$ is identified as 0.3. As $\beta$ increases, the proportion of loss associated with the dynamic construction of directed hypergraphs also escalates. This indicates that $\beta$ is crucial in balancing the trade-off between the complexity of hypergraph construction and the overall performance of the model.
Meanwhile, the threshold $\zeta$ serves as a key parameter in the dynamic construction of the hypergraph, determining which candidate nodes will be included in the hyperedges. The candidate nodes whose correlation with the master node exceeds the threshold $\zeta$ will be incorporated into the hyperedges. An increase in threshold $\zeta$ leads to a decrease in the number of candidate nodes incorporated into the hyperedges. The results show that the optimal value of $\zeta$ is 0.3 for $D_{\textit{Hangzhou}}$ and 0.0 for $D_{\textit{Jinan}}$. This reduction adversely affects the ability to effectively capture the spatiotemporal correlations between intersections, ultimately degrading the performance of ATT. Similar trends are recorded when testing on two synthetic datasets $D_{\textit{4$\times$4}}$ and $D_{\textit{6$\times$6}}$ (We omit the details due to space limitation).
6 Conclusion
@@
In this paper, we propose a multi-agent policy-based framework DHLight that incorporates a directed hypergraph module to capture the correlation in road traffic by dynamically constructing spatio-temporal hyperedges. This innovative approach allows for a deep comprehension of the directed spatio-temporal relationships between multiple intersections and their traffic signals, enabling the obtained model to adapt to varying traffic conditions in real time. Extensive experimental results demonstrate the superiority of DHLight against the baselines in traffic signal control. Our future work is to enhance the robustness of our method by studying the data cleaning schemes to mitigate the impact of missing or noisy existing in the collected data.
{ack}
This work was supported in part by the National Natural Science Foundation of China (Grant No. 62472332) and the Natural Science Foundation of Shanghai (No.25ZR1402511).